# Dozerformer: Sequence Adaptive Sparse Transformer for Multivariate Time Series Forecasting

### v1

@@ Abstract
Transformers have achieved remarkable performance in multivariate time series(MTS) forecasting due to their capability to capture long-term dependencies. However, the canonical attention mechanism has two key limitations: (1) its quadratic time complexity limits the sequence length, and (2) it generates future values from the entire historical sequence. To address this, we propose a Dozer Attention mechanism consisting of three sparse components: (1) Local, each query exclusively attends to keys within a localized window of neighboring time steps. (2) Stride, enables each query to attend to keys at predefined intervals. (3) Vary, allows queries to selectively attend to keys from a subset of the historical sequence. Notably, the size of this subset dynamically expands as forecasting horizons extend. Those three components are designed to capture essential attributes of MTS data, including locality, seasonality, and global temporal dependencies. Additionally, we present the Dozerformer Framework, incorporating the Dozer Attention mechanism for the MTS forecasting task. We evaluated the proposed Dozerformer framework with recent state-of-the-art methods on nine benchmark datasets and confirmed its superior performance. The code will be released after the manuscript is accepted.
@@ 1 Introduction
Multivariate time series (MTS) forecasting endeavors to adeptly model the dynamic evolution of multiple variables over time from their historical records, thereby facilitating the accurate prediction of future values. This task holds significant importance in various applications Petropoulos et al. (2022). The advent of deep learning has significantly advanced MTS forecasting, various methods are proposed based on Recurrent Neural Networks (RNN) Lai et al. (2018), Convolutional Neural Networks (CNN) Shih et al. (2019); Wang et al. (2023). In recent years, the Transformer has demonstrated remarkable efficacy in MTS forecasting. This notable performance can be attributed to its inherent capacity to effectively capture global temporal dependencies across the entirety of the historical sequence. The Transformer is initially proposed for natural language processing(NLP) Vaswani et al. (2017) tasks, with its primary objective of extracting information from words within sentences. It achieved great success in the NLP field and swiftly extended its impact in computer vision Dosovitskiy et al. (2021) and time series analysis Wen et al. (2023). The critical challenge for applying transformers in MTS forecasting is to modify the attention mechanism based on the characteristics of time series data. Informer Zhou et al. (2021), Autoformer Wu et al. (2021), FEDformer Zhou et al. (2022), and Crossformer Zhang & Yan (2023) modified the canonical full attention mechanism to improve its efficiency and achieved promising results in MTS forecasting. Dlinear Zeng et al. (2023) challenges the effectiveness of the transformer by proposing a straightforward linear layer to infer historical records to predictions directly and outperforming aforementioned Transformer-based methods on benchmark datasets.
@@
  • We introduce a sequence-adaptive sparse Dozer Attention mechanism composed of three components: Local, Stride, and Vary. Each component aims to capture temporal dependencies from a select number of effective historical time steps. Notably, the Vary component dynamically expands its utilization of historical time steps as the forecasting horizons extend.
  • We propose the Dozerformer framework for MTS forecasting, incorporating the Dozer Attention mechanism designed to model and predict the seasonal components within MTS data.
  • The experimental results on nine benchmark datasets showcase Dozerformer’s superior performance in terms of both accuracy and efficiency when compared to recent state-of-the-art methods.
@@ 2 Related Work
2.1 MTS Forecasting
MTS forecasting involves predicting values ($X_{pred}\in\mathbb{R}^{O\times D}$) of $D$ variables for future $O$ time steps, given their historical records within the look-back window ($X\in\mathbb{R}^{I\times D}$) spanning $I$ time steps. Deep learning has significantly advanced MTS forecasting. LSTNet Lai et al. (2018) employs CNN and RNN to capture both short-term and long-term temporal dependencies. TPA-LSTM Shih et al. (2019) introduces a temporal pattern attention mechanism to weight hidden states across historical time steps. Graph WaveNet Wu et al. (2019) pioneers the use of Graph Neural Networks (GNN) for MTS forecasting, jointly training an adjacency matrix to model spatial dependencies. MTGNN Wu et al. (2020) is the first GNN-based method for generic MTS forecasting. It utilizes GNN to model correlations among multiple variables and CNN with different filter sizes to model multi-scale temporal dependencies. Transformers Wen et al. (2023) gained popularity in MTS forecasting due to their capacity to capture long-term temporal dependencies. The attention mechanism enables modeling temporal dependencies across all historical time steps. While CNN needs to be stacked, RNN needs to iteratively go through all time steps, and GNN is vulnerable to the adjacency matrix which indicates the correlation between variables. MICN Wang et al. (2023) replaces self-attention with a CNN-based local-global structure to model temporal dependencies across scales efficiently. Dlinear Zeng et al. (2023) challenges transformer effectiveness with a simple linear layer, achieving state-of-the-art performance. PatchTST Nie et al. (2023) highlights the significance of transformers in modeling temporal dependencies at sub-series resolution, independently of variables. In summary, transformers have pushed the boundaries of MTS forecasting.
2.2 Sparse Self Attentions
@@ 3 Method
In this section, we present the Dozerformer method, which incorporates the Dozer attention mechanism. Section 3.1 introduces the Dozerformer framework, which includes a transformer encoder-decoder pair and a linear layer designed for forecasting seasonal and trend components of MTS data, respectively. Section 3.2 provides a comprehensive description of the proposed Dozer attention mechanism, with its primary concept centered around the elimination of query-key pairs between input and output time steps that do not contribute to accuracy.
@@ 3.1 Framework
The MTS forecasting task aims to infer values of $D$ variables for future $O$ time steps $X\in\mathbb{R}^{O\times D}$ from their historical records $X\in\mathbb{R}^{I\times D}$. Figure 2 illustrates the overall framework of Dozerformer. Initially, it decomposes the MTS data into seasonal and trend components, following previous methods Wu et al. (2021); Zhou et al. (2022); Zeng et al. (2023). Subsequently, the transformer and linear models are employed to generate forecasts for the seasonal component $\mathbf{X}_{s}\in\mathbb{R}^{I\times D}$ and the trend component $\mathbf{X}_{t}\in\mathbb{R}^{I\times D}$, respectively.
The dimension invariant (DI) embed Zhang et al. (2023) transforms raw single-channel MTS data into multi-channel feature maps while preserving both the time step and variable dimensions. It further partitions the MTS data into multiple patches along the time step dimension, resulting in patched MTS embeddings denoted as $X_{enc}\in\mathbb{R}^{c\times N_{enc}\times p\times D}$. Here, $c$ represents the number of feature maps, $N_{enc}=\lceil I/p\rceil$ indicates the number of patches for the encoder, and $p$ represents the patch size. Similarly, the decoder’s input, denoted as $X_{dec}\in\mathbb{R}^{c\times N_{dec}\times p\times D}$ (not shown in Figure 2), is derived from $L$ historical time steps and $O$ zero padding. In this context, $N_{dec}=\lceil(L+O)/p\rceil$ indicates the number of patches for the decoder. The transformer model is designed to capture temporal dependencies at multiple time step resolutions (sub-series level). While the transformer encoder and decoder adhere to the canonical transformer architecture Vaswani et al. (2017), they employ the proposed Dozer attention mechanism in place of the canonical full attention. A $1\times 1$ CNN is employed to generate seasonal component predictions based on the learned latent representations from the transformer.
To predict the trend component, we employ a linear layer for the direct inference of trend forecasts from historical trend components. Subsequently, the forecasts for the seasonal and trend components are aggregated by summation to obtain the final predictions denoted as $\mathbf{X}_{pred}\in\mathbb{R}^{O\times D}$.
@@
The canonical scaled dot-product attention is as follows:
$\begin{split}Q,K,V&=\operatorname{Linear}\left(\mathbf{X}_{enc}\right)\\ \operatorname{Attention}\left(Q,K,V\right)&=\operatorname{Softmax}\left(QK^{T}/\sqrt{d_{k}}\right)V\end{split}$ (1)
where the $Q$, $K$, and $V$ represent the query, keys, and values obtained through embedding from the input sequence of the $d$-th series, denoted as $\mathbf{X}{enc}^{d}\in\mathbb{R}^{c\times N{enc}\times p}$. It’s noteworthy that we flatten both the feature map and patch size dimensions of $\mathbf{X}{enc}^{d}$, resulting in $\mathbf{X}{enc}^{d}\in\mathbb{R}^{N_{enc}\times(c\times p)}$, representing the latent representations of patches with a size of $p_{k}$. The canonical transformer exhibits two key limitations: firstly, its encoder’s self-attention is afflicted by quadratic computational time complexity, thereby restricting the length of $N_{enc}$; secondly, its decoder’s cross-attention employs the entire sequence to forecast all future time steps, entailing unnecessary computation for ineffective historical time steps. To overcome these issues, we propose a Dozer attention mechanism, comprising three components: Local, Stride, and Vary.
@@ 3.2.1 Local
$\begin{split}A^{self}_{i,j}&=\begin{cases}q_{i}*k_{j},\qquad&\text{if $j\in\{|i-j|<=\lfloor w/2\rfloor\}$}\\ 0,\qquad&\text{otherwise}\\ \end{cases}\\ A^{cross}_{i,j}&=\begin{cases}q_{i}*k_{j},\qquad&\text{if $j\in\{0<=t-j<=\lfloor w/2\rfloor\}$}\\ 0,\qquad&\text{otherwise}\end{cases}\\ \end{split}$ (2)
In the case of self-attention, the Local component enables each query to selectively attend to a specific set of keys that fall within a defined time window. Specifically, each query in self-attention attends to $\lfloor w/2\rfloor+1$ keys, with $\lfloor w/2\rfloor$ keys representing the neighboring time steps on each side of the query. For cross-attention, the concept of locality differs since future time steps cannot be utilized during forecasting. Instead, the locality is derived from the last $\lfloor w/2\rfloor+1$ time steps, situated at the end of the sequence. These steps encompass the most recent time points up to the moment of forecasting, denoted as time $t$.
@@ 3.2.2 Stride
$A_{i,j}=\begin{cases}q_{i}*k_{j}\qquad&\text{if $j\in\{|i-j|\%interval==0\}$}\\ 0,\qquad&\text{otherwise}\end{cases}$ (3)
@@ 3.2.3 Vary
$A^{cross}_{i,j}=\begin{cases}q_{i}*k_{j},\qquad&\text{if $j\in\{t-1,t-2,\cdots,t-v+i-t\}$ and $i>t$}\\ 0,\qquad&\text{otherwise}\end{cases}$ (4)
@@ 4.1 Experimental settings
Datasets. We conduct experiments on 9 public benchmarks Wu et al. (2021), including 4 ETT datasets(ETTh1, ETTh2, ETTm1, ETTm2), Traffic, Electricity, Weather, Exchange-Rate, and ILI. These are widely used real-world MTS datasets with different characteristics.
@@
Training details. We implemented the proposed Dozerformer in Pytorch and trained with Adam optimizer ($\beta_{1}=0.9$ and $\beta_{2}=0.99$). The initial learning rate is selected from $\{5e-5,1e-4,5e-4,1e-3\}$ via grid search and updated using the Cosine Annealing scheme. The transformer encoder layer is set to 2 and the decoder layer is set to 1. We select the look-back window size of historical time steps utilized from $\{96,192,336,720\}$ via grid search except for the ILI dataset which is set to $120$. The patch size is selected from $\{24,48,96\}$ via grid search. The main results are repeated six times using seeds $1,2022,2023,2024,2025$, and $2026$. The ablations study and parameter sensitivity results are obtained from random seed $1$.

### v2

# Sparse Transformer with Local and Seasonal Adaptation for Multivariate Time Series Forecasting

@@ Abstract
Transformers have achieved remarkable performance in multivariate time series(MTS) forecasting due to their capability to capture long-term dependencies. However, the canonical attention mechanism has two key limitations: (1) its quadratic time complexity limits the sequence length, and (2) it generates future values from the entire historical sequence. To address this, we propose a Dozer Attention mechanism consisting of three sparse components: (1) Local, each query exclusively attends to keys within a localized window of neighboring time steps. (2) Stride, enables each query to attend to keys at predefined intervals. (3) Vary, allows queries to selectively attend to keys from a subset of the historical sequence. Notably, the size of this subset dynamically expands as forecasting horizons extend. Those three components are designed to capture essential attributes of MTS data, including locality, seasonality, and global temporal dependencies. Additionally, we present the Dozerformer Framework, incorporating the Dozer Attention mechanism for the MTS forecasting task. We evaluated the proposed Dozerformer framework with recent state-of-the-art methods on nine benchmark datasets and confirmed its superior performance. The experimental results indicate that excluding a subset of historical time steps from the time series forecasting process does not compromise accuracy while significantly improving efficiency. Code is available at https://github.com/GRYGY1215/Dozerformer.
@@ 1 Introduction
Multivariate time series (MTS) forecasting endeavors to adeptly model the dynamic evolution of multiple variables over time from their historical records, thereby facilitating the accurate prediction of future values. This task holds significant importance in various applications Petropoulos et al. (2022). The advent of deep learning has significantly advanced MTS forecasting, various methods are proposed based on Recurrent Neural Networks (RNN) Lai et al. (2018b); Zhang et al. (2024b; 2021), Convolutional Neural Networks (CNN) Shih et al. (2019); Wang et al. (2023). In recent years, the Transformer has demonstrated remarkable efficacy in MTS forecasting. This notable performance can be attributed to its inherent capacity to effectively capture global temporal dependencies across the entirety of the historical sequence. The Transformer is initially proposed for natural language processing(NLP) Vaswani et al. (2017) tasks, with its primary objective of extracting information from words within sentences. It achieved great success in the NLP field and swiftly extended its impact in computer vision Dosovitskiy et al. (2021) and time series analysis Wen et al. (2023). The critical challenge for applying transformers in MTS forecasting is to modify the attention mechanism based on the characteristics of time series data. Informer Zhou et al. (2021) introduced a ProbSparse attention mechanism that restricts each key to interact only with a fixed number of the most significant queries. This model employs a generative-style decoder to produce predictions for multiple time steps in a single forward pass, aiming to capture temporal dependencies at the individual time step level. Autoformer Wu et al. (2021) proposed an Auto-Correlation mechanism to model temporal dependencies at the sub-series level. It also introduced a time series decomposition block that utilizes moving averages to separate seasonal and trend components from raw MTS data. FEDformer Zhou et al. (2022) utilizes a Frequency Enhanced Block and Frequency Enhanced Attention to discern patterns in the frequency domain. It achieves linear computational complexity by selectively sampling a limited number of Fourier components for the attention mechanism’s queries, keys, and values. Pyraformer Liu et al. (2022) introduces a pyramidal graph organized as a tree structure where each parent node has several child nodes. It aims to capture data patterns at multiple scales by enabling each parent node to merge the sub-series from its child nodes. As a result, the scale of the graph exponentially increases based on the number of child nodes. Crossformer Zhang & Yan (2023) introduced a Two-Stage Attention mechanism that applies full attention across both time steps and variable dimensions. It also proposed a three-level framework designed to capture hierarchical latent representations of MTS data by merging tokens produced at lower levels. Dlinear Zeng et al. (2023) challenges the effectiveness of the transformer by proposing a straightforward linear layer to infer historical records to predictions directly and outperforming aforementioned Transformer-based methods on benchmark datasets. MICN Wang et al. (2023) introduced a CNN-based framework that achieves linear computational complexity. It features a local-global structure designed to capture long-term temporal dependencies in MTS data, surpassing the performance of both the full attention and Auto-Correlation mechanisms Wu et al. (2021). Notably, it incorporates a multi-scale hybrid decomposition block that uses various kernel sizes to decompose MTS data into seasonal and trend components effectively. PatchTST Nie et al. (2023) utilizes a full attention mechanism to model the temporal dependencies of individual variables at the sub-series level. It proposes a patching procedure that divides the input sequence into multiple time-step patches, significantly reducing computational complexity. Additionally, it adopts self-supervised learning to generate latent representations of MTS data
@@
  • We introduce a sequence-adaptive sparse Dozer Attention mechanism composed of three components: Local, Stride, and Vary. Each component aims to capture temporal dependencies from a select number of effective historical time steps. Notably, the Vary component dynamically expands its utilization of historical time steps as the forecasting horizons extend. To the best of our knowledge, Dozer Attention is the first mechanism that adaptively utilizes historical time steps based on the forecasting horizons. It addresses the limitation of existing attention mechanisms by overcoming their quadratic computational complexity, thereby enhancing efficiency.
  • We propose the Dozerformer framework for MTS forecasting, incorporating the Dozer Attention mechanism designed to model and predict the seasonal components within MTS data. It comprises a transformer encoder-decoder pair coupled with a linear layer, yielding a straightforward yet effective architecture. The Dozerformer framework achieved a slight improvement in accuracy and a significant improvement in efficiency.
  • The experimental results on nine benchmark datasets showcase Dozerformer’s superior performance in terms of both accuracy and efficiency when compared to recent state-of-the-art methods. Additionally, our extensive ablation studies demonstrate the effectiveness and robustness of the proposed Dozer Attention mechanism.
@@ 2 Related Work
MTS Forecasting
MTS forecasting involves predicting values ($X_{pred}\in\mathbb{R}^{O\times D}$) of $D$ variables for future $O$ time steps, given their historical records within the look-back window ($X\in\mathbb{R}^{I\times D}$) spanning $I$ time steps. Deep learning has significantly advanced MTS forecasting. LSTNet Lai et al. (2018b) employs CNN and RNN to capture both short-term and long-term temporal dependencies. TPA-LSTM Shih et al. (2019) introduces a temporal pattern attention mechanism to weight hidden states across historical time steps. Graph WaveNet Wu et al. (2019) pioneers the use of Graph Neural Networks (GNN) for MTS forecasting, jointly training an adjacency matrix to model spatial dependencies. MTGNN Wu et al. (2020) is the first GNN-based method for generic MTS forecasting. It utilizes GNN to model correlations among multiple variables and CNN with different filter sizes to model multi-scale temporal dependencies. Transformers Wen et al. (2023) gained popularity in MTS forecasting due to their capacity to capture long-term temporal dependencies. The attention mechanism enables modeling temporal dependencies across all historical time steps. While CNN needs to be stacked, RNN needs to iteratively go through all time steps, and GNN is vulnerable to the adjacency matrix which indicates the correlation between variables. MICN Wang et al. (2023) replaces self-attention with a CNN-based local-global structure to model temporal dependencies across scales efficiently. Dlinear Zeng et al. (2023) challenges transformer effectiveness with a simple linear layer, achieving state-of-the-art performance. PatchTST Nie et al. (2023) highlights the significance of transformers in modeling temporal dependencies at sub-series resolution, independently of variables. In summary, transformers have pushed the boundaries of MTS forecasting.
Sparse Self Attentions
@@ 3 Method
This section presents the Dozerformer method, incorporating the Dozer attention mechanism. We first introduce the Dozerformer framework, comprising a transformer encoder-decoder pair and a linear layer designed to forecast the seasonal and trend components of MTS data. Subsequently, we provide a comprehensive description of the proposed Dozer attention mechanism, focusing on eliminating query-key pairs between input and output time steps that do not contribute to accuracy.
@@ 3.1 Framework
The MTS forecasting task aims to infer values of $D$ variables for future $O$ time steps $X\in\mathbb{R}^{O\times D}$ from their historical records $X\in\mathbb{R}^{I\times D}$. Figure 2 illustrates the overall framework of Dozerformer. Initially, it decomposes the MTS data into seasonal and trend components, following previous methods Wu et al. (2021); Zhou et al. (2022); Zeng et al. (2023); Wang et al. (2023). Subsequently, the transformer and linear models are employed to generate forecasts for the seasonal component $\mathbf{X}_{s}\in\mathbb{R}^{I\times D}$ and the trend component $\mathbf{X}_{t}\in\mathbb{R}^{I\times D}$, respectively.
@@
The dimension invariant (DI) embed Zhang et al. (2024a) transforms raw single-channel MTS data into multi-channel feature maps while preserving both the time step and variable dimensions. It further partitions the MTS data into multiple patches along the time step dimension, resulting in patched MTS embeddings denoted as $X_{enc}\in\mathbb{R}^{c\times N_{enc}\times p\times D}$. Here, $c$ represents the number of feature maps, $N_{enc}=\lceil I/p\rceil$ indicates the number of patches for the encoder, and $p$ represents the patch size. Similarly, the decoder’s input, denoted as $X_{dec}\in\mathbb{R}^{c\times N_{dec}\times p\times D}$ (not shown in Figure 2), is derived from $L$ historical time steps and $O$ zero padding. In this context, $N_{dec}=\lceil(L+O)/p\rceil$ indicates the number of patches for the decoder. The transformer model is designed to capture temporal dependencies at multiple time step resolutions (sub-series level). While the transformer encoder and decoder adhere to the canonical transformer architecture Vaswani et al. (2017), they employ the proposed Dozer attention mechanism in place of the canonical full attention. A $1\times 1$ CNN is employed to generate seasonal component predictions based on the learned latent representations from the transformer.
To predict the trend component, we employ a linear layer for the direct inference of trend forecasts from historical trend components. Subsequently, the forecasts for the seasonal and trend components are aggregated by summation to obtain the final predictions denoted as $\mathbf{X}_{pred}\in\mathbb{R}^{O\times D}$. A detailed step-by-step computation procedure is introduced in Supplementary Section 2.
@@ 3.2 Dozer Attention
The canonical scaled dot-product attention is as follows:
$\begin{split}Q,K,V&=\operatorname{Linear}\left(\mathbf{X}_{enc}^{d}\right)\\ \operatorname{Attention}\left(Q,K,V\right)&=\operatorname{Softmax}\left(QK^{T}/\sqrt{d_{k}}\right)V\end{split}$ (1)
where the $Q$, $K$, and $V$ represent the queries, keys, and values obtained through embedding from the input sequence of the $d$-th series, denoted as $\mathbf{X}_{enc}^{d}\in\mathbb{R}^{c\times N_{enc}\times p}$. It’s noteworthy that we flatten both the feature map and patch size dimensions of $\mathbf{X}_{enc}^{d}$, resulting in $\mathbf{X}_{enc}^{d}\in\mathbb{R}^{N_{enc}\times(c\times p)}$, representing the latent representations of patches with a size of $p$. The scaling factor $d_{k}$ denotes the dimension of the vectors in queries and keys. The canonical transformer exhibits two key limitations: firstly, its encoder’s self-attention is afflicted by quadratic computational time complexity, thereby restricting the length of $N_{enc}$; secondly, its decoder’s cross-attention employs the entire sequence to forecast all future time steps, entailing unnecessary computation for ineffective historical time steps. To overcome these issues, we propose a Dozer Attention mechanism, comprising three components: Local, Stride, and Vary.
@@ 3.2.1 Local
$\begin{split}A^{self}_{i,j}&=\begin{cases}q_{i}*k_{j},\qquad&\text{if $j\in\{|i-j|\leq\lfloor w/2\rfloor\}$}\\ 0,\qquad&\text{otherwise}\\ \end{cases}\\ A^{cross}_{i,j}&=\begin{cases}q_{i}*k_{j},\qquad&\text{if $j\in\{0\leq t-j\leq\lfloor w/2\rfloor\}$}\\ 0,\qquad&\text{otherwise}\end{cases}\\ \end{split}$ (2)
Where $A$ represents the attention matrix that records the values of production between queries and keys. The superscripts $self$ and $cross$ denote self-attention and cross-attention, respectively. The subscripts $i$ and $j$ represent the locations or time steps of the query vector and key vector along the temporal dimension, respectively. $t$ specifies the time step at which forecasting is conducted, and $w$ represents the window size for the Local component. In the case of self-attention, the Local component enables each query to selectively attend to a specific set of keys that fall within a defined time window. Specifically, each query in self-attention attends to $12\times\lfloor w/2\rfloor+$ 1 keys, with $\lfloor w/2\rfloor$ keys representing the neighboring time steps on each side of the query. For cross-attention, the concept of locality differs since future time steps cannot be utilized during forecasting. Instead, the locality is derived from the last $\lfloor w/2\rfloor+1$ time steps, situated at the end of the sequence. These steps encompass the most recent time points up to the moment of forecasting, denoted as time $t$.
@@ 3.2.2 Stride
$A_{i,j}=\begin{cases}q_{i}*k_{j},\qquad&\text{if $j\in\{|i-j|\mod s=0\}$}\\ 0,\qquad&\text{otherwise}\end{cases}$ (3)
@@ 3.2.3 Vary
$A^{cross}_{i,j}=\begin{cases}q_{i}*k_{j},\qquad&\text{if $j\in\{|i-j|<v+i-t$\}and $i>t$}\\ 0,\qquad&\text{otherwise}\end{cases}$ (4)
@@ 4.1 Experimental settings
Datasets. We conduct experiments on 9 public benchmarks Wu et al. (2021), including 4 ETT datasets(ETTh1, ETTh2, ETTm1, ETTm2), Traffic, Electricity, Weather, Exchange-Rate, and ILI. These are widely used real-world MTS datasets with different characteristics. Details are introduced in Supplementary Section 1.
@@
Training details. We implemented the proposed Dozerformer in Pytorch and trained with Adam optimizer ($\beta_{1}=0.9$ and $\beta_{2}=0.99$). The initial learning rate is selected from $\{5e-5,1e-4,5e-4,1e-3\}$ via grid search and updated using the Cosine Annealing scheme. The transformer encoder layer is set to 2 and the decoder layer is set to 1. We select the look-back window size of historical time steps utilized from $\{96,192,336,720\}$ via grid search except for the ILI dataset which is set to $120$. The patch size is selected from $\{24,48,96\}$ via grid search. The main results, summarized in Table 1, are average values derived from six repetitions using seeds $1,2022,2023,2024,2025$, and $2026$. The ablations study and parameter sensitivity results are obtained from random seed $1$.